### v1

@@ Abstract
The China Space Station Telescope (CSST) is a next-generation Stage-IV sky survey telescope, distinguished by its large field of view (FoV), high image quality, and multi-band observation capabilities. It can simultaneously conduct precise measurements of the Universe by performing multi-color photometric imaging and slitless spectroscopic surveys. The CSST is equipped with five scientific instruments, i.e. Multi-band Imaging and Slitless Spectroscopy Survey Camera (SC), Multi-Channel Imager (MCI), Integral Field Spectrograph (IFS), Cool Planet Imaging Coronagraph (CPI-C), and THz Spectrometer (TS). Using these instruments, the CSST is expected to make significant contributions and discoveries across various astronomical fields, including cosmology, galaxy and active galactic nuclei (AGN), the Milky Way and nearby galaxies, stars, exoplanets, Solar System objects, astrometry, and transients and variable sources. This review aims to provide a comprehensive overview of the CSST instruments, observational capabilities, data products, and scientific potential.
@@
dark energy, dark matter, cosmological constraint
@@ 1 Introduction
As a Stage-IV survey telescope, the China Space Station Telescope (CSST) is an off-axis three-mirror anastigmat (TMA) telescope with a primary mirror of 2 meters in diameter, designed to probe the Universe with extremely high levels of precision [12, 13, 14, 15, 16]. The CSST is scheduled to be launched in 2027, which will be co-orbiting with the China Space Station (CSS or Tiangong) at an altitude of about 400 kilometers, and it can rendezvous and dock with the CSS for on-orbit servicing (OOS). Aiming at different scientific goals and astronomical targets, it is equipped with five scientific instruments, i.e. Multi-band Imaging and Slitless Spectroscopy Survey Camera (SC), Multi-Channel Imager (MCI), Integral Field Spectrograph (IFS), Cool Planet Imaging Coronagraph (CPI-C), and THz Spectrometer (TS). The design parameters of the telescope and scientific instruments are summarized in Table 1.
@@
In this review, we briefly introduce the CSST instruments, observation strategy, data processing, and discuss its scientific potential. The paper is organized as follows. In Section 2, we list the scientific objectives of the CSST. In Section 3, the design of the survey platform and primary optical system of the telescope and the five scientific instruments are introduced. In Section 4, we discuss the CSST observations for different instruments. In Section 5, we introduce the CSST data processing and data release. In Section 6, the details of the CSST science are discussed. We summarize in Section 7.
@@ 3.3 Survey camera
The focal plane contains 2.6 billion pixels, corresponding to an angular size of $0.074^{\prime\prime}$ per pixel, a FoV of $\sim$1.1 deg2 and dynamic image quality $R_{\rm EE80}<0.15^{\prime\prime}$. The layout of the main focal plane is shown in Figure 2. The main focal plane is divided into 7 photometric imaging bands (i.e. $NUV$, $u$, $g$, $r$, $i$, $z$, and $y$) and 3 slitless spectroscopic bands ($GU$, $GV$, and $GI$, with an average spectral resolution of $R\geq 200$), covering the wavelength range of 255-1000 nm. The transmission curves of the seven photometric bands and three slitless spectroscopic bands are shown in Figure 3. The photometric imaging bands are located in the central area, and the slitless spectroscopic bands are arranged in the upper and lower “L”-shaped areas. The charge-coupled devices (CCDs) of the main focal plane are fixedly installed with the filters or gratings.
@@
Figure 3: The transmission curves of the seven photometric imaging bands, i.e. $NUV$, $u$, $g$, $r$, $i$, $z$, and $y$ (solid curves), and three slitless spectroscopic bands, i.e. $GU$, $GV$, and $GI$ (dashed curves) for the CSST-SC [13].
@@ 3.4.1 MCI
The MCI is equipped with a total of 30 filters, with 10 filters for each channel. Among them, 4 filters cover the entire FoV, and 6 filters cover 1/4 of the FoV. The near-UV channel $CBU$ filter, optical-blue $CBV$ filter, and optical-red $CBI$ filter basically cover the spectral range of the channel, and will be used in the observation of the MCI XDF of $\sim$300 arcmin2, with magnitude limits $CBV\geq 30$, and $CBU$ and $CBI\geq 29$. The filters $NUV$, $u$, $r$, $z$ and $y$ can be used to assist the flux calibration of the SC. The remaining ones include the narrow-band, mid-band and broad-band filters, mainly used in scientific research such as high-redshift (high-$z$) and neighboring galaxies, and ionized gas [18, 19].
@@ 4.3 MCI observation
The MCI XDF observations will be conducted in the $CBU$, $CBV$ and $CBI$ bands in the Earth’s shadow region, with a cumulative exposure of 510,000 s in each band and magnitude limits of no less than 29, 30, and 29, respectively. Six directions will be selected for the XDFs, totaling 300 arcmin2, of which four directions will be used to observe galaxy clusters. The HST XDF is tentatively selected as one of the CSST XDFs, and is planned to be completed in the first two years of scientific operation. The remaining XDFs will be completed roughly simultaneously in the following seven years. MCI can conduct observations in parallel with IFS or CPI-C, obtaining additional valuable observational data.
@@ 4.6 TS observation
    THz spectral line survey: Single-point deep observation of high and low-mass star-forming regions, carbon stars, and low-mass protostars. The cumulative effective observation time for a single source is expected to be $50-100$ hours.
@@ 5.1 Data processing
The CSST Scientific Data Processing System generates level-1 and level-2 data products based on the original observational data (level-0 data). Level-1 data refers to scientific data that has been corrected for instrument effects and calibrated, and level-2 data refers to data products that have been further analyzed on the basis of level-1 data to form catalogs and complete the extraction of preliminary scientifically usable parameters, including model photometry of galaxies, photometric redshift (photo-$z$), stellar atmospheric parameters, etc. The CSST data products include:
@@ 5.2 Data management
In addition, the Digital Telescope System is also under development, providing digital support for understanding the CSST operating status and abnormal monitoring. To support CSST scientific data research, the Space Application System plans to establish a CSST Data Science Center and connect to the public cloud system. On the one hand, it can provide the storage and computing resources for CSST long-term tasks, and on the other hand, it can make full use of the artificial intelligence (AI) technology.
@@ 6 Science
The CSST Science Committee is organized to guide CSST scientific research and review top-level scientific policies. Four science centers have been established, i.e the Peking University Science Center, the NAOC (National Astronomical Observatories, Chinese Academy of Sciences) Science Center, the Yangtze River Delta Region Science Center, and the Guangdong-Hong Kong-Macau Greater Bay Area Science Center, to provide support platforms for the CSST scientific teams. The NAOC Science Center also serves as the CSST Joint Science Center, collaborating with the science centers to jointly plan and manage CSST scientific projects.
@@ 6.1.1 Dark energy
The physical property of dark energy is mainly determined by the ratio of its pressure $p$ to energy density $\rho$, i.e. the equation of state $w\equiv p/\rho$. For example, in the $\Lambda$CDM model, dark energy is the cosmological constant and can be seen as vacuum energy, that we have $w=-1$. On the other hand, in dynamical dark energy models, $w$ is a function of the redshift $z$ and can evolve with time, which can be seen as a scalar field, e.g. quintessence with $w>-1$ [20], phantom with $w<-1$ [21], and quintom with $w$ crossing $-1$ [22]. Therefore, it is crucial to reconstruct the evolution history of $w$ from astronomical observations for studying the nature of dark energy.
@@
CSST can explore the nature of dark energy by various cosmological probes, e.g. weak and strong gravitational lensing, galaxy angular and redshift-space clustering, galaxy cluster, cosmic void, SN Ia, BAO, redshift-space distortion (RSD), etc.
@@ 6.1.1.1   Weak gravitational lensing
Figure 5: The predicted contour maps of $\Omega_{\rm m}$ vs. $\sigma_{8}$ and $w_{0}$ vs. $w_{a}$ derived from the CSST 3$\times$2pt, galaxy spec-$z$, and galaxy cluster surveys. The 68.3% and 95.5% CLs are shown. The gray dotted lines show the parameter fiducial values. The systematics have been considered in the constraint process, such as the shear calibration, photo-$z$ uncertainties, intrinsic alignment, galaxy bias, non-linear and instrumental effects.
@@
On the other hand, for the next generation of weak gravitational lensing surveys like CSST, the required photo-$z$ accuracy is extremely high, with the accuracies of the redshift bias $\Delta z$ being at the level of one thousandth and the scatter $\sigma_{z}$ is at the level of one percent. CSST has the wavelength coverage from near-UV to near-IR, with a total of seven bands, which has advantages for accurately measuring photo-$z$. Using the traditional spectral energy distribution (SED) fitting method, it is found that $\sigma_{z}$ can achieve $\sim 0.03$, and the outlier or fraction of catastrophic redshift is less than $5\%$ [14, 18] for the CSST photometric survey. When adopting the machine learning methods with appropriate training data from spectroscopic surveys, $\sigma_{z}$ can be better than $0.02$ and the outlier fraction will be less than $3\%$ [26, 27, 28, 29, 30, 31]. When considering the measurements in the other bands from other telescopes, e.g. $Euclid$ $Y$, $J$ and $H$ bands, the photo-$z$ accuracy can be further significantly improved [14, 32].
@@
It is expected that the CSST 17,500 deg2 wide-field photometric survey can obtain more than one billion galaxy images with a spatial resolution higher than $0.15^{\prime\prime}$ and a number density of $\sim 28$ arcmin-2, and the galaxy redshift distribution has a peak abound $z=0.6-0.7$ and can extend to $z>4$ [15]. Based on the results derived from Jiutian simulations [33] and theoretical predictions, the shear two-point correlation functions or power spectra in different tomographic redshift bins can be precisely measured at $30<z<$ 3, and the constraint accuracy of $w$ can reach $\sim 10\%$ level when only using the data of the CSST shear measurements [34, 35, 36, 37].
@@
In addition, since the weak lensing effect is closely related to the LSS, the main signal comes from the nonlinear structures, which is highly non-Gaussian. Hence, the shear two-point correlation cannot reflect all the cosmological information. The peak statistics of weak lensing are concentrated in nonlinear regions and highly complementary to the shear two-point statistics [38, 39, 40]. The constraint strength of $w$ can be further improved when cooperating the two weak lensing probes.
The weak lensing survey also can cross-correlate with galaxy angular clustering measurement to obtain three two-point correlation functions (3$\times$2pt) for cosmological constraints, i.e. shear, galaxy angular and galaxy-galaxy lensing correlation functions or power spectra. The 3$\times$2pt data can effectively break the degeneracies between model parameters, and improve the constraints on the cosmological parameters and significantly suppress the effects of systematics, e.g. photo-$z$ uncertainties [36]. In Figure 5, we show the constraint results of $\Omega_{\rm m}-\sigma_{8}$ and $w_{0}-w_{a}$ from the CSST 3$\times$2pt surveys (see blue contours), where $\Omega_{\rm m}$ is total matter density parameter and $\sigma_{8}$ is the amplitude of matter fluctuation on the scale of $8\ h^{-1}{\rm Mpc}$. The systematics from the shear calibration (multiplicative and additive errors), photo-z uncertainties (redshift bias and scatter), intrinsic alignment, galaxy bias, non-linear effects and instruments, have been considered in the analysis. The details can be found in [35]. We can find that the CSST 3$\times$2pt surveys can constrain $\Omega_{\rm m}$, $\sigma_{8}$, and equation of state of dark energy with accuracies $\sim 3\%$, $\sim 1\%$, and $\sim 8\%$, respectively.
@@ 6.1.1.2   Galaxy redshift-space clustering
Besides multi-band photometric imaging survey, CSST also will simultaneously perform the slitless spectroscopic wide-field survey with three bands, i.e. $GU$, $GV$, and $GI$. The spectral resolution $R$ is higher than 200, and the spatial resolution is about $0.3^{\prime\prime}$. It is expected that the CSST slitless spectroscopic wide-field survey can obtain more than one hundred million galaxy spectra with a surface number density $2-3$ arcmin-2 in 17,500 deg2 survey area. The galaxy redshift distribution has a peak around $z=0.3$ and can extend to $z>1.5$ derived from the zCOSMOS catalog, which means that the galaxy number density $n_{\rm g}$ is greater than $10^{-2}\,h^{3}{\rm Mpc^{-3}}$ at $0.60<z<$ 0.6 and $10^{-3}\,h^{3}{\rm Mpc^{-3}}$ at $1.20.6<z<$ 1.2 [15].
In a spectroscopic survey, the galaxy clustering is measured in redshift space, i.e. the redshift-space galaxy power spectrum $P^{i}_{\rm g}(k,\mu)$ in a spectroscopic redshift (spec-$z$) bin $i$, where $k$ is the wavenumber and $\mu$ is the cosine of the angle between $k$ and the line of sight. Theoretically, $P_{\rm g}(k,\mu)$ is dependent on several effects, e.g. the Alcock-Paczynski (AP) effect [41], Kaiser effect [42], damping effect, and shot noise. The uncertainty of $P_{\rm g}(k,\mu)$ is tightly related to the galaxy number density and the spec-$z$ accuracy.
@@
In Figure 5, the contour maps of $\Omega_{\rm m}-\sigma_{8}$ and $w_{0}-w_{a}$ from the CSST slitless spectroscopic galaxy survey have been shown. We can find that the constraint accuracies of $\Omega_{\rm m}$, $\sigma_{8}$, and equation of state of dark energy are about $10\%$, $2\%$, and $20\%$, respectively, which are weaker than the results of the CSST 3$\times$2pt surveys.
@@ 6.1.1.3   Galaxy cluster
Galaxy cluster is the largest virialized system in the Universe with a typical mass $10^{14}-10^{15}\ M_{\odot}$, where the inward pull of gravity is balanced by the outward motion of the galaxies to form a gravitationally bound system that has reached a state of equilibrium. Since the formation of galaxy clusters is tightly related to the LSS, they can be adopted to constrain the property of dark energy, e.g. using the number counts of galaxy clusters.
Accurately measuring the redshift and mass of galaxy clusters is crucial in the cosmological study of galaxy clusters. CSST can simultaneously conduct photometric imaging and slitless spectroscopic surveys, which is highly advantageous for the identification and verification of galaxy clusters, as well as for measuring their redshift and mass. We note that the redshift and members of a cluster can be determined by the CSST spectroscopic observation, and the mass can be derived from the gravitational lensing and richness measurements in CSST photometric survey. Therefore, in principle, we can precisely perform the number counts of galaxy clusters in a mass range at a given redshift for the CSST surveys.
At $1.50<z<$ 1.5, which is the redshift range covered by both CSST photometric and spectroscopic surveys, about 200,000-400,000 galaxy clusters can be detected according to the theoretical estimation based on the halo mass function, and the peak of cluster redshift distribution is around $z=0.6$ [35, 44]. The constraint results of $\Omega_{\rm m}-\sigma_{8}$ and $w_{0}-w_{a}$ from the CSST number counts of galaxy clusters are shown in red contours in Figure 5. We can find that the constraint accuracies of $\Omega_{\rm m}$, $\sigma_{8}$, and equation of state of dark energy are $\sim 5\%$, $\sim 1\%$, and $\sim 13\%$, respectively, which are better than that from the CSST galaxy redshift-space clustering measurements. However, we should notice that, for other parameters, e.g. $\Omega_{\rm b}$ and $n_{\rm s}$, the constraint power of the CSST number counts of galaxy clusters could be worse than the galaxy redshift-space clustering, due to low sensitivity to these parameters [35].

### v3

@@ Abstract
The Chinese Space Station Survey Telescope (CSST) is an upcoming Stage-IV sky survey telescope, distinguished by its large field of view (FoV), high image quality, and multi-band observation capabilities. It can simultaneously conduct precise measurements of the Universe by performing multi-color photometric imaging and slitless spectroscopic surveys. The CSST is equipped with five scientific instruments, i.e. Multi-band Imaging and Slitless Spectroscopy Survey Camera (SC), Multi-Channel Imager (MCI), Integral Field Spectrograph (IFS), Cool Planet Imaging Coronagraph (CPI-C), and THz Spectrometer (TS). Using these instruments, CSST is expected to make significant contributions and discoveries across various astronomical fields, including cosmology, galaxies and active galactic nuclei (AGN), the Milky Way and nearby galaxies, stars, exoplanets, Solar System objects, astrometry, and transients and variable sources. This review aims to provide a comprehensive overview of the CSST instruments, observational capabilities, data products, and scientific potential.
@@
Telescope, cosmology, galaxy
@@ 1 Introduction
As a Stage-IV survey telescope, the Chinese Space Station Survey Telescope (CSST) is an off-axis three-mirror anastigmat (TMA) telescope with a primary mirror of 2 meters in diameter, designed to probe the Universe with extremely high levels of precision [12, 13, 14, 15, 16]. The CSST is scheduled to be launched in 2027, which will be co-orbiting with the China Space Station (CSS or Tiangong) at an altitude of about 400 kilometers, and it can rendezvous and dock with the CSS for on-orbit servicing (OOS). Aiming at different scientific goals and astronomical targets, it is equipped with five scientific instruments, i.e. Multi-band Imaging and Slitless Spectroscopy Survey Camera (SC), Multi-Channel Imager (MCI), Integral Field Spectrograph (IFS), Cool Planet Imaging Coronagraph (CPI-C), and THz Spectrometer (TS). The design parameters of the telescope and scientific instruments are summarized in Table 1.
@@
In this review, we briefly introduce the CSST instruments, observation strategy, data processing, and discuss its scientific potential. The paper is organized as follows. In Section 2, we list the scientific objectives of the CSST. In Section 3, the design of the survey platform, primary optical system, and the five scientific instruments are introduced. In Section 4, we discuss the CSST observations for different instruments. In Section 5, we introduce the CSST data processing and data release. In Section 6, the details of the CSST science are discussed. We summarize in Section 7.
@@ 3.3 Survey camera
The focal plane contains 2.6 billion pixels, corresponding to an angular size of $0.074^{\prime\prime}$ per pixel, a FoV of $\sim$1.1 deg2 and dynamic image quality $R_{\rm EE80}<0.15^{\prime\prime}$. The layout of the main focal plane is shown in Figure 2. The main focal plane is divided into 7 photometric imaging bands (i.e. $NUV$, $u$, $g$, $r$, $i$, $z$, and $y$) and 3 slitless spectroscopic bands ($GU$, GVGV, and $GI$, with an average spectral resolution of $R\geq 200$), covering the wavelength range of 255-1000 nm. The transmission curves of the seven photometric bands and three slitless spectroscopic bands are shown in Figure 3. The photometric imaging bands are located in the central area, and the slitless spectroscopic bands are arranged in the upper and lower “L”-shaped areas. The charge-coupled devices (CCDs) of the main focal plane are fixedly installed with the filters or gratings.
@@
Figure 3: The transmission curves of the seven photometric imaging bands, i.e. $NUV$, $u$, $g$, $r$, $i$, $z$, and $y$ (solid curves), and three slitless spectroscopic bands, i.e. $GU$, GVGV, and GIGI (dashed curves) for the CSST-SC [13].
@@ 3.4.1 MCI
The MCI is equipped with a total of 30 filters, with 10 filters for each channel. Among them, 4 filters cover the entire FoV, and 6 filters cover 1/4 of the FoV. The near-UV channel $CBU$ filter, optical-blue $CBV$ filter, and optical-red $CBI$ filter basically cover the spectral range of the channel, and will be used in the observation of the MCI XDF of $\sim$300 arcmin2, with magnitude limits $CBV\geq 30$, and $CBU$ and $CBI\geq 29$. The filters $NUV$, $u$, $r$, $z$ and $y$ can be used to assist the flux calibration of the SC. The remaining ones include the narrow-band, mid-band and broad-band filters, mainly used in scientific research such as high-redshift (high-$z$) and neighboring galaxies, and ionized gas [17, 18, 19].
@@ 4.3 MCI observation
The MCI XDF observations will be conducted in the $CBU$, CBVCBV and $CBI$ bands in the Earth’s shadow region, with a cumulative exposure of 510,000 s in each band and magnitude limits of no less than 29, 30, and 29, respectively. Six directions will be selected for the XDFs, totaling 300 arcmin2, of which four directions will be used to observe galaxy clusters. The HST XDF is tentatively selected as one of the CSST XDFs, and is planned to be completed in the first two years of scientific operation. The remaining XDFs will be completed roughly simultaneously in the following seven years. MCI can conduct observations in parallel with IFS or CPI-C, obtaining additional valuable observational data.
@@ 4.6 TS observation
    THz spectral line survey: Single-point deep observation of high and low-mass star-forming regions, carbon stars, and low-mass protostars. The cumulative effective observation time for a single source is expected to be 5010050-100 hours.
@@ 5.1 Data processing
The CSST Scientific Data Processing System generates level-1 and level-2 data products based on the original observational data (level-0 data). Level-1 data refers to scientific data that has been corrected for instrument effects and calibrated. Level-2 data refers to data products that have been further analyzed on the basis of level-1 data to form catalogs and complete the extraction of preliminary scientifically usable parameters, including model photometry of galaxies, photometric redshift (photo-$z$), stellar atmospheric parameters, etc. The CSST data products include:
@@ 5.2 Data management
In addition, the Digital Telescope System is also under development, providing digital support for understanding the CSST operating status and abnormal monitoring. To support CSST scientific data research, the CSST Data Science Center is established, which connects to the public cloud system. On the one hand, it can provide the storage and computing resources for CSST long-term tasks, and on the other hand, it can make full use of the artificial intelligence (AI) technology.
@@ 6 Science
The CSST Science Committee is organized to guide CSST scientific research and review top-level scientific policies. Five science centers have been established, i.e the Peking University Science Center, the NAOC (National Astronomical Observatories, Chinese Academy of Sciences) Science Center, the Yangtze River Delta Region Science Center, the Guangdong-Hong Kong-Macau Greater Bay Area Science Center, and CSST Data Science Center, to provide support platforms for the CSST scientific teams. The NAOC Science Center also serves as the CSST Joint Science Center, collaborating with all the science centers to jointly plan and manage CSST scientific projects.
@@ 6.1.1 Dark energy
[…]
@@
[…]
CSST can explore the nature of dark energy with high precision by various cosmological probes in a large redshift range, e.g. weak and strong gravitational lensing, galaxy angular and redshift-space clustering power spectra, galaxy cluster, cosmic void, SN Ia, BAO, redshift-space distortion (RSD), etc.
@@ 6.1.1.1 Weak gravitational lensing
[…]
@@
On the other hand, for the upcoming weak gravitational lensing surveys like CSST, the required photo-$z$ accuracy is extremely high, with the accuracies of the redshift bias $\Delta z$ being at the level of one thousandth and the scatter $\sigma_{z}$ is at the level of one percent. CSST has the wavelength coverage from near-UV to near-IR, with a total of seven bands, which has advantages for accurately measuring photo-$z$. Using the traditional spectral energy distribution (SED) fitting method, it is found that $\sigma_{z}$ can achieve $\sim 0.03$, and the outlier or fraction of catastrophic redshift is less than $5\%$ [14, 18] for the CSST photometric survey. When adopting the machine learning methods with appropriate training data from spectroscopic surveys, $\sigma_{z}$ can be better than $0.02$ and the outlier fraction will be less than $3\%$ [39, 40, 41, 42, 43, 44]. When considering the measurements in the other bands from other telescopes, e.g. EuclidEuclid $Y$, $J$ and $H$ bands, the photo-$z$ accuracy can be further significantly improved [14, 45].
@@
It is expected that the CSST 17,500 deg2 wide-field photometric survey can obtain more than one billion galaxy images with a spatial resolution higher than $0.15^{\prime\prime}$ and a number density of $\sim 28$ arcmin-2, and the galaxy redshift distribution has a peak abound z=0.60.7z=0.6-0.7 and can extend to $z>4$ [15]. Based on the results derived from Jiutian simulations [46] and theoretical predictions, the shear two-point correlation functions or power spectra in different tomographic redshift bins can be precisely measured at $0<z<3$, and the constraint accuracy of $w$ can reach $\sim 10\%$ level when only using the data of the CSST shear measurements [47, 48, 49, 50].
@@
In addition, since the weak lensing effect is closely related to the LSS, the main signal comes from the nonlinear structures, which is highly non-Gaussian. Hence, the shear two-point correlation cannot reflect all the cosmological information. On the other hand, the peak statistics of weak lensing are concentrated in nonlinear regions and highly complementary to the shear two-point statistics [51, 52, 53]. The constraint strength of $w$ can be further improved when cooperating the two weak lensing probes.
[…]
@@ 6.1.1.2 Galaxy redshift-space clustering
[…]
In a spectroscopic survey, the galaxy clustering is measured in redshift space, i.e. the redshift-space galaxy power spectrum $P^{i}_{\rm g}(k,\mu)$ at the linear regime in a spectroscopic redshift (spec-$z$) bin $i$, where $k$ is the wavenumber and $\mu$ is the cosine of the angle between $k$ and the line of sight. Theoretically, $P_{\rm g}(k,\mu)$ is dependent on several effects, e.g. the Alcock-Paczynski (AP) effect [54], Kaiser effect [55], damping effect, and shot noise. The uncertainty of $P_{\rm g}(k,\mu)$ is tightly related to the galaxy selection function, survey volume, galaxy number density and the spec-$z$ accuracy. The spec-$z$ accuracy is one of the most important sample-selection criteria in the CSST cosmological study.
@@
[…]
@@ 6.1.1.3 Galaxy cluster
Galaxy cluster is the largest virialized system in the Universe with a typical mass $10^{14}-10^{15}\ M_{\odot}$, which is formed as a gravitationally bound system that has reached a state of equilibrium. Since the formation of galaxy clusters is tightly related to the LSS, they can be adopted to constrain the property of dark energy, e.g. using the number counts of galaxy clusters.
Accurately measuring the redshift and mass of galaxy clusters is crucial in the cosmological study of galaxy clusters. CSST can simultaneously conduct photometric imaging and slitless spectroscopic surveys, which is highly advantageous for the identification and verification of galaxy clusters, as well as for measuring their redshift and mass. We note that the redshift and members of a cluster can be determined by the CSST spectroscopic observation, and the mass can be derived from the gravitational lensing and richness measurements in CSST photometric survey. Therefore, in principle, we can precisely perform the number counts of galaxy clusters in a mass range ($\gtrsim 10^{14}\ M_{\odot}$) at a given redshift for the CSST surveys.
At $0<z<1.5$, which is the redshift range covered by both CSST photometric and spectroscopic surveys, about 200,000-400,000 galaxy clusters can be detected according to the theoretical estimation based on the halo mass function, and the peak of cluster redshift distribution is around $z=0.6$ [48, 57]. The constraint results of $\Omega_{\rm m}-\sigma_{8}$ and $w_{0}-w_{a}$ from the CSST number counts of galaxy clusters are shown in red contours in Figure 5. We can find that the constraint accuracies of $\Omega_{\rm m}$, $\sigma_{8}$, and equation of state of dark energy are $\sim 5\%$, $\sim 1\%$, and $\sim 13\%$, respectively, which are better than that from the CSST galaxy redshift-space clustering measurements. However, we should notice that, for other parameters, e.g. $\Omega_{\rm b}$ and $n_{\rm s}$, the constraint power of the CSST number counts of galaxy clusters could be worse than the galaxy redshift-space clustering, due to low sensitivity to these parameters [48].